# Modelling and Classifying the Components of a Literature Review

### v1

@@ Abstract
Previous work has demonstrated that AI methods for analysing scientific literature benefit significantly from annotating sentences in papers according to their rhetorical roles, such as research gaps, results, limitations, extensions of existing methodologies, and others. Such representations also have the potential to support the development of a new generation of systems capable of producing high-quality literature reviews. However, achieving this goal requires the definition of a relevant annotation schema and effective strategies for large-scale annotation of the literature. This paper addresses these challenges by 1) introducing a novel annotation schema specifically designed to support literature review generation and 2) conducting a comprehensive evaluation of a wide range of state-of-the-art large language models (LLMs) in classifying rhetorical roles according to this schema. To this end, we also present Sci-Sentence, a novel multidisciplinary benchmark comprising 700 sentences manually annotated by domain experts and 2,240 sentences automatically labelled using LLMs. We evaluate 37 LLMs on this benchmark, spanning diverse model families and sizes, using both zero-shot learning and fine-tuning approaches. The experiments yield several novel insights that advance the state of the art in this challenging domain. First, the current generation of LLMs performs remarkably well on this task when fine-tuned on high-quality data, achieving performance levels above 96% F1. Second, while large proprietary models like GPT-4o achieve the best results, some lightweight open-source alternatives also demonstrate excellent performance. Finally, enriching the training data with semi-synthetic examples generated by LLMs proves beneficial, enabling small encoders to achieve robust results and significantly enhancing the performance of several open decoder models.
keywords:
Literature Review , Systematic Literature Review , Artificial Intelligence , Large Language Models , Text Classification
journal: Engineering Applications of Artificial Intelligence
[…]
[label1]organization=Knowledge Media Institute, The Open University, addressline=Walton Hall, city=Milton Keynes, postcode=MK7 6AA, country=UK
[…]
[label2]organization=Department of Business and Law, University of Milano Bicocca, addressline=Piazza dell’Ateneo Nuovo, 1, city=Milan, postcode=20126, country=IT
@@ 1 Introduction
A literature review or related work section is a fundamental component of a research paper, as it provides the necessary background, highlights the research gap, and justifies the research objectives. It also serves to summarise relevant literature in educational settings, aiding students and researchers in understanding the state of the art regarding a certain topic. However, crafting a high-quality literature review remains a challenging task, even for experienced researchers. It requires comprehensive knowledge of the relevant literature, which is increasingly difficult to maintain due to the growing volume of published research [1] and the continual need for updates to ensure relevance [2]. In addition, it demands the ability to synthesise this information into a clear and structured discussion that highlights key research directions, theoretical frameworks, and open challenges in the field.
The artificial intelligence (AI) and natural language processing (NLP) communities have been actively researching the analysis and automatic generation of related work sections for more than 15 years. The latter task has traditionally been framed as related work summarization [3] and typically involves three steps: identifying documents relevant to an input query, understanding the relationships and interactions among these documents, and producing a coherent summary [3]. The landscape of related work generation has shifted significantly with the advent of large language models (LLMs), which are capable of producing fluent, natural-sounding summaries of research papers. This advancement has led to the emergence of various systems that claim to generate scientific reports and even to compose full scientific papers by leveraging and referencing relevant scholarly literature [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. These systems typically follow a retrieval-augmented generation (RAG) pipeline: a user query (e.g., a request for a summary on a specific topic) is used to retrieve pertinent papers from a vector database, and the retrieved documents are then provided as context to the LLM for response generation [16]. However, despite the grammatical fluency and surface coherence of the outputs produced by current LLM-based approaches, the quality of the resulting literature reviews remains very limited. This is primarily because such outputs tend to consist of uncritical summaries of individual papers, rather than structured and analytical reviews [17]. In contrast, prior research has shown that literature reviews in academic writing are expected to follow a coherent structure and fulfil specific rhetorical functions [18]. These include identifying research gaps, highlighting methodological or conceptual limitations, synthesising findings across studies, discussing theoretical and practical implications, and proposing directions for future research.
@@
To enable the development of a new generation of systems capable of producing high-quality literature reviews, we argue that it is essential to begin with a more sophisticated representation of the claims made in relevant papers, which would characterise each sentence according to its rhetorical role. For instance, this would enable a system to retrieve all sentences that discuss research challenges within a specific area and generate a focused overview based on future research directions derived from that content. Indeed, previous work [19, 20, 21, 22] has shown that AI methods for analyzing scientific literature benefit significantly from annotating sentences in papers according to their rhetorical roles, such as research gaps, results, and limitations. Texts annotated with such roles have been shown to facilitate the analysis of the evolution of scientific knowledge [19], as well as to assist in identifying [20, 21] and predicting [22] the significance of scientific concepts and contributions.
The intuition behind employing a richer representation of the literature, in which each sentence is associated with its rhetorical role, to support systems for literature review generation raises two crucial research questions. First, what type of annotation schema can effectively assist both systems and users in the task of generating literature reviews? Second, can current NLP technologies, particularly LLMs, be leveraged to accurately identify the rhetorical roles of sentences within research papers?
@@
This paper addresses these challenges by 1) introducing a novel annotation schema specifically designed to support literature review generation, and 2) performing a comprehensive evaluation of a wide range of state-of-the-art LLMs in classifying rhetorical roles according to this schema.
We began by developing an annotation schema inspired by prior studies on rhetorical structure [23, 18], which categorises scientific sentences into seven classes: Overall, Research Gap, Description, Result, Limitation, Extension, and Other. Based on this schema, we created a new publicly available resource, Sci-Sentence, a multidisciplinary benchmark that includes 700 sentences manually annotated by domain experts and 2,240 sentences automatically labelled using LLMs. We then evaluated 37 LLMs spanning various model families and sizes on this benchmark, using both zero-shot learning and fine-tuning approaches.
@@
These experiments yielded several novel insights that advance the state of the art in this challenging domain. First, the current generation of LLMs performs remarkably well on this task when fine-tuned on high-quality datasets such as Sci-Sentence, reaching performance levels above 96% F1. Second, while large proprietary models like GPT-4o achieve the best results, lightweight open-source alternatives, such as SuperNova-Medius and Nemotron-8B, also demonstrate excellent performance. Third, although decoder-only models achieve the highest overall scores, small and scalable encoder-based models pre-trained on domain-relevant data, such as SciBERT, also achieve solid performance. Therefore, they represent a practical solution for efficiently processing large volumes of text. Finally, enriching the training data with semi-synthetic examples generated by LLMs has proven beneficial. This approach enables small encoders to achieve robust results and significantly enhances the performance of several open decoder models.
@@
The remainder of this paper is structured as follows. Section 2 reviews related work, including established literature review frameworks and existing approaches for classifying sentences in scientific articles. Section 3 defines the task, describes the development of the annotation schema, and presents the novel benchmarks. Section 4 details the experimental methodology and describes the models and approaches used to classify scientific sentences. Section 5 presents the experimental results, and Section 6 provides additional analysis of the performance of different methods, optimization techniques, and the effectiveness of semi-synthetic data. Finally, Section 7 concludes the paper and outlines potential directions for future research.
@@ 2.1 Frameworks based on the Rhetorical Structure of Scientific Papers
Rhetorical Structure Theory is a framework for text organisation that has inspired applications in discourse analysis, text generation, psycholinguistics, and computational linguistics [24]. It has also been extensively applied to the study, understanding, and generation of scientific and scholarly texts. In particular, three main types of analyses have emerged from its application to scientific and academic literature: genre analysis [25, 26], zoning analysis [27, 28, 29, 30], and discourse analysis [31, 32].
@@
Genre analysis is commonly employed in the field of linguistics due to its effectiveness in manual rhetorical analysis and its pedagogical value in academic writing instruction [25, 33, 34, 35]. It provides a structured guide to analysing and creating introductions of scientific papers or review chapters of doctoral theses. In contrast, zoning and discourse analyses are studied in NLP research, as they provide machine-interpretable categories and enable fine-grained, sentence-level annotation throughout the entire scholarly document. While zoning analysis focuses on identifying the rhetorical function of individual sentences within the scientific argument [27, 29], discourse analysis examines the textual coherence, meaning, and structural relationships at the sentence level [23].
@@
In the domain of related work summarization [3], only a limited number of studies have investigated genre or discourse analysis. Wang et al. [36], inspired by the genre analysis, proposed the CaRS model, which describes how academic writers structure introductions by establishing, justifying, and presenting their work. The authors also introduced RSGen, a transformer-based model that employs a two-step decoding process: first creating a rhetorical plan, then generating the related work content. However, RSGen achieves only moderate performance and suffers from issues such as error propagation in classification and limited generalizability to alternative rhetorical schemas. Khoo et al. [23], drawing on the discourse-based analysis, focused instead on the foundational task of understanding human-written literature reviews through manual analysis. They examined the macro-level discourse structure of literature reviews in information science journals, developing a coding schema with 12 categories. Although they identified distinct structural and content-related differences between integrative and descriptive literature reviews, their coding schema faced challenges, including difficulties in differentiating categories, occasional low inter-rater reliability for specific categories, and a lack of operationalization in automated computational methods.
@@
In this study, we adopt a discourse-based approach inspired by the findings of Jaidka, Khoo & Na [18], which emphasize the critical role of sentences in the generation of related work section and recognizing the lack of a well-defined structure in existing research [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. In particular, we seek to enhance the coding schema introduced by Khoo et al. [23], with the objective of formulating machine-interpretable categories that facilitate their automation by AI systems.
@@ 2.2 Approaches for Classifying Sentences in Related Work Section Generation
The literature presents a variety of approaches for classifying sentences in scholarly articles. Several methods focus on the rhetorical structure of the paper [23], most often targeting the discourse around citations, including citation function [37, 38, 39, 40, 41, 42], citation intent [43, 44], and citation sentiment [45, 46, 47]. Other approaches aim to associate sentences or text segments with relevant topics [48, 49], often selected from one of the many knowledge organisation systems used to categorise scientific literature [50]. Another category of systems focuses on extracting research entities (e.g., tasks, methods, materials) linked by semantic relations [51]. The richer representations of the literature produced by these systems are often encoded in knowledge graphs [52] (e.g., SemOpenAlex [53], ORKG [54], AI-KG [55], CS-KG [56], Nano-publications [57]) and have been shown to support effectively scientometric analyses [58, 59], intelligent systems for exploring the literature [60], and, increasingly, conversational systems [61] and question-answering methods [62, 63, 64]. However, as noted by several recent studies [65, 66, 17], the output of existing approaches does not adequately support the generation of related work sections. Therefore, in this paper, we propose a novel sentence classification approach explicitly designed to support related work analysis and generation.
@@
To the best of our knowledge, only two approaches for related work generation incorporate sentence classification into their pipeline [67, 68], and both specifically focus on characterising citations. Before discussing these methods, it is useful to first introduce the strategies employed in related work summarization. These strategies are commonly classified as either extractive or abstractive. Extractive systems utilise sentences as their units of analysis, producing a paragraph by concatenating selected sentences in a specific order [3, 69, 70, 71, 72, 73]. The paragraph does not have any division and lacks the structure of a human-written literature review [66]. In contrast, abstractive systems process input such as excerpts or paragraphs generating fluent paragraphs [74, 67, 68, 75, 76, 77, 78, 79, 80, 81, 82, 6, 7]. However, their outputs often exhibit deficiencies, such as the absence of transitional sentences, improper citation ordering, or as in extractive approaches the lack of structure [66].
@@
The two approaches that use sentence classification are both abstractive methods. Xing et al. [67] trained a BERT [83] model to classify sentences as explicit citations, which directly name the source, or implicit citations, which refer to the work without naming it. Ge et al. [68] fine-tuned SciBERT [84] to categorise citation sentences as positive, negative, or neutral, depending on whether they emphasise contributions, highlight shortcomings, or provide objective descriptions of the cited work.

### v2

@@ Abstract
Previous work has demonstrated that AI methods for analysing scientific literature benefit significantly from annotating sentences in papers according to their rhetorical roles, such as research gaps, results, limitations, extensions of existing methodologies, and others. Such representations also have the potential to support the development of a new generation of systems capable of producing high-quality literature reviews. However, achieving this goal requires the definition of a relevant annotation schema and effective strategies for large-scale annotation of the literature. This paper addresses these challenges in two ways: 1) it introduces a novel, unambiguous annotation schema that is explicitly designed for reliable automatic processing, and 2) it presents a comprehensive evaluation of a wide range of large language models (LLMs) on the task of classifying rhetorical roles according to this schema. To this end, we also present Sci-Sentence, a novel multidisciplinary benchmark comprising 700 sentences manually annotated by domain experts and 2,240 sentences automatically labelled using LLMs. We evaluate 37 LLMs on this benchmark, spanning diverse model families and sizes, using both zero-shot learning and fine-tuning approaches. The experiments reveal that modern LLMs achieve strong results on this task when fine-tuned on high-quality data, surpassing 96% F1, with both large proprietary models such as GPT-4o and lightweight open-source alternatives performing well. Moreover, augmenting the training set with semi-synthetic LLM-generated examples further boosts performance, enabling small encoders to achieve robust results and substantially improving several open decoder models.
@@ 1 Introduction
A literature review or related work section is a fundamental component of a research paper, as it provides the necessary background, highlights the research gap, and justifies the research objectives (Booth et al., , 2021). It also serves to summarise relevant literature in educational settings, aiding students and researchers in understanding the state of the art regarding a certain topic. However, crafting a high-quality literature review remains a challenging task, even for experienced researchers. It requires comprehensive knowledge of the relevant literature, which is increasingly difficult to maintain due to the growing volume of published research (Bornmann and Mutz, , 2015) and the continual need for updates to ensure relevance (Moher et al., , 2007). In addition, it demands the ability to synthesise this information into a clear and structured discussion that highlights key research directions, theoretical frameworks, and open challenges in the field.
@@
In recent years, artificial intelligence (AI) and natural language processing (NLP) research has focused on analysis and automatic generation of scientific literature. Previous studies (Wang et al., , 2022; Groza, , 2013; Widyantoro et al., , 2013; Teufel and Moens, , 2002) have shown that AI methods for analyzing scientific literature benefit significantly from annotating sentences in papers according to their rhetorical roles, such as research gaps, results, and limitations. Texts annotated with such roles have been demonstrated to facilitate the analysis of the evolution of scientific knowledge (Wang et al., , 2022) and to assist in identifying (Groza, , 2013; Widyantoro et al., , 2013) and predicting (Teufel and Moens, , 2002) the significance of scientific concepts and contributions.
More recently, the rise of large language models (LLMs) has led to the emergence of various systems that claim to generate scientific analysis by leveraging relevant scholarly literature (Shi et al., , 2023; Zhang et al., 2024c, ; Li and Ouyang, 2024a, ; Martin-Boyle et al., , 2024; Zhang et al., 2024a, ; Nishimura et al., , 2024; Li et al., 2024d, ; Ma et al., , 2025; Agarwal et al., , 2025; Liu et al., , 2025; Beger and Henneking, , 2025; Wang et al., , 2025). Many of these systems focus on the specific task of related work generation, which involves producing coherent and informative text that synthesises and discusses relevant aspects of the existing literature, similar to the content typically found in the literature review or related work sections of scientific papers. However, despite the surface coherence of the outputs produced by current LLM-based approaches, the quality of the resulting literature reviews remains limited. This limitation primarily arises because these outputs often consist of uncritical summaries of individual papers rather than structured and analytical reviews that take into consideration important concepts such as research gaps and limitations (Bolanos et al., , 2024). To support the development of systems capable of producing high-quality literature reviews, we argue that it is essential to integrate rhetorical frameworks for literature analysis which identify the functional role of each sentence with the natural language generation capabilities of LLMs. Such integration would enable systems to leverage a richer representation of the input text, allowing, for example, the retrieval of all sentences that discuss research gaps within a specific domain in order to generate coherent overviews focused on future research directions.
The first step toward achieving this integration is to address the task of classifying text according to its rhetorical structure, thereby producing a representation that can support downstream systems for literature analysis and evaluation. In this study, we focus particularly on recent transformer-based architectures, which have demonstrated strong performance in related tasks involving the classification of technical text (Fields et al., , 2024). This objective gives rise to two main research questions. First, what kind of annotation schema can enable a fully automatic characterisation of sentences in literature reviews? Second, how can current NLP technologies, particularly transformer models, reliably identify the rhetorical roles of sentences in research papers?
@@
This paper tackles these challenges by: 1) introducing a novel annotation schema explicitly designed to support reliable automatic processing, and 2) performing a comprehensive evaluation of state-of-the-art transformer models on a human-curated benchmark constructed according to this schema.
As first step, we developed an annotation schema inspired by prior studies on rhetorical structure (Khoo et al., , 2011; Jaidka et al., , 2013), which categorises scientific sentences into seven classes: Overall, Research Gap, Description, Result, Limitation, Extension, and Other. Unlike earlier approaches (Khoo et al., , 2011), which were primarily developed for manual annotation and used a few potentially ambiguous labels requiring contextual interpretation beyond the capability of automated classifiers, our schema defines clear and unambiguous categories that can be consistently annotated by machine learning models.
Based on this schema, we created a new publicly available resource, Sci-Sentence, a multidisciplinary benchmark that includes 700 sentences manually annotated by domain experts and 2,240 sentences automatically labelled using LLMs. We then evaluated several categories of transformer models on this dataset, covering encoder-only, encoder-decoder, and decoder-only architectures. The evaluation compared models of different sizes (1–3B, 7–8B, and large-scale LLMs), various prominent model families (e.g., BERT, T5, Llama, Mistral, Gemma, Phi), and both proprietary and open-source variants. In total, we tested 37 models, organised into distinct categories, using both zero-shot and fine-tuning settings. The aim was to assess the overall performance of transformer-based models and to gain deeper insights into the most effective approaches and optimisation strategies for this challenging task. We also conducted an in-depth error analysis to identify the categories that pose the greatest challenges for automatic recognition across different model types. For example, we found that the Limitation and Description categories are particularly difficult, as they can sometimes appear in forms that may be misconstrued as other types. Furthermore, we compared two fine-tuning optimisation techniques, LoRA and NEFT, and observed that LoRA achieved superior results for large decoder models. Finally, we investigated the impact of training on synthetic data across different model architectures.
@@
The experiments yielded several novel insights that advance the state of the art in this challenging domain. First, the current generation of LLMs performs remarkably well on this task when fine-tuned on high-quality datasets such as Sci-Sentence, reaching performance levels above 96% F1. Second, while large proprietary models like GPT-4o achieve the best results, lightweight open-source alternatives, such as SuperNova-Medius and Nemotron-8B, also demonstrate excellent performance. Third, although decoder-only models achieve the highest overall scores, small and scalable encoder-based models pre-trained on domain-relevant data, such as SciBERT, also achieve solid performance. Therefore, they represent a practical solution for efficiently processing large volumes of text. Finally, enriching the training data with semi-synthetic examples generated by LLMs has proven very beneficial. This solution enables small encoders to achieve robust results and significantly enhances the performance of several open decoder models.
@@
The remainder of this paper is structured as follows. Section 2 reviews related work, including established literature review frameworks and existing approaches for classifying sentences in scientific articles. Section 3 defines the task, describes the development of the annotation schema, and presents the novel benchmarks. Section 4 details the experimental methodology and describes the models and approaches used to classify scientific sentences. Section 5 presents the experimental results, with additional analysis on the performance of different methods, optimization techniques, and the effectiveness of semi-synthetic data. Finally, Section 6 concludes the paper and outlines potential directions for future research.
@@ 2.1 Frameworks based on the Rhetorical Structure of Scientific Papers
Rhetorical Structure Theory is a framework for text organisation that has inspired applications in discourse analysis, text generation, psycholinguistics, and computational linguistics (Taboada and Mann, , 2006). It has also been extensively applied to the study, understanding, and generation of scientific and scholarly texts. In particular, three main types of analyses have emerged from its application to scientific and academic literature: genre analysis (Swales and Swales, , 1990; Kanoksilapatham, , 2005), zoning analysis (Teufel et al., , 1999; Mizuta et al., , 2006; Teufel et al., , 2009; Liakata et al., , 2010), and discourse analysis (Mann and Thompson, , 1988; Prasad et al., , 2008).
@@
Genre analysis is commonly employed in the field of linguistics due to its effectiveness in manual rhetorical analysis and its pedagogical value in academic writing instruction (Swales and Swales, , 1990; Swales et al., , 2004; Kwan, , 2006; Bastola and Ho, , 2023). It provides a structured guide to analysing and creating introductions of scientific papers or review chapters of doctoral theses. In contrast, zoning and discourse analyses are studied in NLP research, as they provide machine-interpretable categories and enable fine-grained, sentence-level annotation throughout the entire scholarly document. While zoning analysis focuses on identifying the rhetorical function of individual sentences within the scientific argument (Teufel et al., , 1999, 2009), discourse analysis examines the textual coherence, meaning, and structural relationships at the sentence level (Khoo et al., , 2011).
@@
In the domain of related work generationonly a limited number of studies have investigated genre or discourse analysis. Wang et al., 2024b , inspired by the genre analysis, proposed the CaRS model, which describes how academic writers structure introductions by establishing, justifying, and presenting their work. The authors also introduced RSGen, a transformer-based model that employs a two-step decoding process: first creating a rhetorical plan, then generating the related work content. However, RSGen achieves only moderate performance and suffers from issues such as error propagation in classification and limited generalizability to alternative rhetorical schemas. Khoo et al., (2011), drawing on the discourse-based analysis, focused instead on the foundational task of understanding human-written literature reviews through manual analysis. They examined the macro-level discourse structure of literature reviews in information science journals, developing a coding schema with 12 categories. Nevertheless, the coding schema presents several challenges that hinder its suitability for fully automated adoption. First, some categories are defined ambiguously and can be easily misinterpreted even by human annotators, often requiring contextual knowledge that an automated classifier would not possess. Second, certain categories are overly broad, which may lead to confusion and reduced consistency in downstream systems. Finally, the schema was not originally designed for computational processing and therefore lacks reliable training datasets and clear, machine-readable classification guidelines.
@@
In this study, we adopt a discourse-based approach inspired by the findings of Jaidka, Khoo & Na (Jaidka et al., , 2013), which emphasizes the critical role of sentences in the generation of related work section and recognizes the lack of a well-defined structure in existing research (Shi et al., , 2023; Zhang et al., 2024c, ; Li and Ouyang, 2024a, ; Martin-Boyle et al., , 2024; Zhang et al., 2024a, ; Nishimura et al., , 2024; Li et al., 2024d, ; Ma et al., , 2025; Agarwal et al., , 2025; Liu et al., , 2025; Beger and Henneking, , 2025; Wang et al., , 2025). In particular, we seek to enhance the coding schema introduced by Khoo et al., (2011), with the objective of formulating machine-interpretable categories that facilitate their automation by AI systems.
@@ 2.2 Approaches for Classifying Sentences in Related Work Section Generation
The literature presents a variety of approaches for classifying sentences in scholarly articles. Several methods focus on the rhetorical structure of the paper (Khoo et al., , 2011), most often targeting the discourse around citations, including citation function (Garfield et al., , 1965; Dong and Schäfer, , 2011; Teufel et al., , 2006; Jurgens et al., , 2018; Tuarob et al., , 2019; Zhao et al., , 2019), citation intent (Cohan et al., , 2019; Lauscher et al., , 2021), and citation sentiment (Athar, , 2011; Athar and Teufel, , 2012; Ravi et al., , 2018). Other approaches aim to associate sentences or text segments with relevant topics (Salatino et al., , 2022; Masoumi and Khajavi, , 2023), often selected from one of the many knowledge organisation systems used to categorise scientific literature (Salatino et al., , 2025). Another category of systems focuses on extracting research entities (e.g., tasks, methods, materials) linked by semantic relations (Dessí et al., 2022a, ). The richer representations of the literature produced by these systems are often encoded in knowledge graphs (Peng et al., , 2023) (e.g., SemOpenAlex (Färber et al., , 2023), ORKG (Jaradeh et al., , 2019), AI-KG (Dessì et al., , 2020), CS-KG (Dessí et al., 2022b, ), Nano-publications (Kuhn et al., , 2016)) and have been shown to support effectively scientometric analyses (Salatino et al., , 2020, 2023), intelligent systems for exploring the literature (Angioni et al., , 2021), and, increasingly, conversational systems (Meloni et al., , 2023) and question-answering methods (Auer et al., , 2023; Lehmann et al., , 2024; Banerjee et al., , 2023). However, as noted by several recent studies (Li and Ouyang, , 2022; Li and Ouyang, 2024b, ; Bolanos et al., , 2024), the output of existing approaches does not adequately support the generation of related work sections. Therefore, in this paper, we propose a novel sentence classification approach explicitly designed to support related work analysis and generation.
@@
To the best of our knowledge, only two approaches for related work generation incorporate sentence classification into their pipeline (Xing et al., , 2020; Ge et al., , 2021), and both specifically focus on characterising citations. Before discussing these methods, it is useful to first introduce the strategies employed in related work generation. These strategies are commonly classified as either extractive or abstractive. Extractive systems utilise sentences as their units of analysis, producing a paragraph by concatenating selected sentences in a specific order (Hoang and Kan, , 2010; Hu and Wan, , 2014; Wang et al., 2019b, ; Chen and Zhuge, , 2019; Wang et al., 2019a, ; Deng et al., , 2021). The paragraph does not have any division and lacks the structure of a human-written literature review (Li and Ouyang, 2024b, ). In contrast, abstractive systems process input such as excerpts or paragraphs generating fluent paragraphs (AbuRa’ed et al., , 2020; Xing et al., , 2020; Ge et al., , 2021; Chen et al., , 2021; Luu et al., , 2020; Jung et al., , 2022; Li et al., 2022a, ; Chen et al., , 2022; Li et al., 2023a, ; Liu et al., , 2023; Gu and Hahnloser, , 2024; Li and Ouyang, 2024a, ; Martin-Boyle et al., , 2024). However, their outputs often exhibit deficiencies, such as the absence of transitional sentences, improper citation ordering, or as in extractive approaches the lack of structure (Li and Ouyang, 2024b, ).
@@
The two approaches that use sentence classification are both abstractive methods. Xing et al., (2020) trained a BERT (Kenton and Toutanova, , 2019) model to classify sentences as explicit citations, which directly name the source, or implicit citations, which refer to the work without naming it. Ge et al., (2021) fine-tuned SciBERT (Beltagy et al., , 2019) to categorise citation sentences as positive, negative, or neutral, depending on whether they emphasise contributions, highlight shortcomings, or provide objective descriptions of the cited work.